# From Flat to Feeling: A Feasibility and Impact Study on Dynamic Facial Emotions in AI-Generated Avatars

### v1

@@ 4.1 Comparison of Audio and Visual Versus Visual-Only Conditions
Variable t-statistic p-value
Mean
Audio+Visual
Mean
Visual-Only
Facial expressions 2.21 0.03 3.29 3.86
Visual appearance 1.39 0.17 3.40 3.80
Voice tone N/A N/A 4.0 N/A
Dialogue content N/A N/A 3.86 N/A
Empathy level 0.75 0.46 2.83 3.06
Ultimately, these results highlight a critical trade-off: while auditory signals are indispensable for effectively conveying anger and resolving ambiguities in high-arousal states, their inclusion risks exposing subtle audiovisual mismatches that can undermine perceived realism. Achieving seamless synchronization between vocal prosody and facial kinematics thus emerges as a prerequisite for maximizing both emotional clarity and immersion in virtual interactions.
4.2 Comparison of Emory Versus Amelia
@@
Audio+Visual Visual-Only
Variable t-statistic p-value Mean_Amelia Mean_Emory t-statistic p-value Mean_Amelia Mean_Emory
Facial exp. 3.496 0.001 3.343 2.800 3.824 0.001 4.000 3.429
Angry 0.000 1.000 2.571 2.571 1.759 0.088 2.314 1.829
Sad 2.451 0.020 4.114 3.629 2.481 0.018 4.086 3.343
Joy -1.139 0.263 1.086 1.171 -1.604 0.118 1.200 1.429
Fear 1.186 0.244 2.086 1.857 0.502 0.619 2.314 2.229
Sad Disgust 2.604 0.014 2.428 1.971 -1.099 0.280 1.800 1.971
Facial exp. 3.421 0.002 3.543 2.914 0.560 0.579 3.629 3.514
Angry -3.446 0.002 3.114 3.829 2.926 0.006 2.171 1.429
Sad 2.758 0.009 3.714 3.257 1.625 0.113 2.371 2.000
Joy 0.572 0.571 1.143 1.086 -1.429 0.162 1.829 2.143
Fear 0.517 0.609 2.000 1.914 0.709 0.483 1.857 1.714
Angry Disgust -2.666 0.012 2.114 2.743 3.053 0.004 1.942 1.486
Facial exp. 1.995 0.054 3.486 3.171 2.119 0.041 3.914 3.514
Angry -1.435 0.160 1.000 1.057 -2.946 0.006 1.086 1.400
Sad -1.000 0.324 1.057 1.114 -3.340 0.002 1.143 1.743
Joy 2.445 0.020 4.200 3.771 6.065 0.000 4.371 2.914
Fear -1.160 0.254 1.114 1.229 -2.321 0.026 1.143 1.571
Joy Disgust -1.675 0.103 1.057 1.171 -2.927 0.006 1.229 1.743
A comparative analysis of the emotional expressiveness specific to each avatar revealed distinct perceptual differences between Amelia, the female-presenting avatar, and Emory, the male-presenting avatar (Table 4). Participants consistently assigned higher ratings to the realism of Amelia’s facial expressions across both audio+visual and visual-only conditions. Statistically significant differences (highlighted in gray in Table 4) emerged in four of the six emotion-condition pairings: sadness in both the audio+visual ($p=0.001$) and visual-only ($p=0.001$) conditions, anger in the audio+visual condition ($p=0.002$), and joy in the visual-only condition ($p=0.041$). These findings indicate a consistent perceptual preference for Amelia’s facial animations across various emotional contexts and presentation modalities, a preference likely influenced by her morphological design.
One potential explanation for Amelia’s superior expressiveness lies in her softer facial geometry and the subtlety of her micro-expressions, which appeared to align more closely with viewers’ expectations of emotional authenticity. For instance, her rounded jawline and gentle brow dynamics effectively amplified the perceived realism of sadness (e.g., downturned lips) and joy (e.g., symmetrical smiles). This observation resonates with previous research on anthropomorphic design, which suggests a correlation between such features and perceived warmth [37].
However, this pattern diverged in the context of anger recognition, where Emory demonstrated superior performance ($p=0.002$). Amelia’s more subdued brow contraction and reduced lip tension appeared to attenuate the visual cues typically associated with aggression. Consequently, in the visual-only condition, 43% of participants misclassified Amelia’s expression of anger as sadness. This finding is consistent with appraisal theory, which underscores the importance of congruent visual and auditory signals for disambiguating high-arousal emotions [38, 39]. In contrast, Emory’s angular facial features and more pronounced brow kinematics served to amplify the perception of aggression. This highlights a potential design trade-off: while a subtler morphology may enhance the realism of low-arousal emotions, it risks ambiguity when portraying high-arousal states. This trade-off is also reflected in findings from multimodal deep learning frameworks, which indicate that the joint modeling of auditory and visual cues improves emotion recognition accuracy, particularly for high-arousal emotions like anger [40]. Furthermore, these perceptual divergences align with cross-cultural research on the interpretation of facial expressions, which demonstrates that morphological features and subtle dynamic expressions can significantly influence emotion attribution. Such research shows that specific facial movements may be interpreted differently depending on both cultural background and structural facial cues [41].
@@ 4.3 Qualitative Feedback
Participants provided reflective qualitative feedback on what stood out the most about the avatars. These reflections revealed distinct and important observations regarding emotional expressiveness, realism, and immersion, providing directions for enhancing avatar technologies in interactive simulations.
@@ 5 Discussion
This research successfully demonstrates the technical viability of a distributed architecture that integrates Unreal Engine 5’s MetaHuman and NVIDIA Omniverse Audio2Face for generating real-time, high-fidelity emotional expressions in AI-driven child avatars. Our empirical evaluation yielded several critical insights. Firstly, the system can synthesize recognizable emotional states, with low-arousal emotions like sadness and joy achieving high identification rates irrespective of whether audio was present. Secondly, the accurate recognition of high-arousal emotions, notably anger, is significantly dependent on congruent vocal cues, with performance dropping substantially in visual-only conditions. Thirdly, avatar morphology plays a crucial role in emotion perception; the softer-featured "Amelia" avatar was generally perceived as more lifelike, whereas the more angular "Emory" avatar more effectively conveyed anger, highlighting a potential trade-off between perceived warmth and the unambiguous display of high-arousal states.
A persistent and significant challenge underscored by this study is audiovisual desynchrony. The marked increase in perceived facial realism when audio was removed suggests that even subtle temporal or qualitative mismatches between auditory and visual channels can considerably disrupt the user’s sense of realism. This phenomenon, also noted in participant feedback regarding lip-sync issues and expressions not matching dialogue, remains a key design bottleneck. Achieving seamless synchronization is not merely a technical refinement but a prerequisite for maximizing both emotional clarity and immersion. This difficulty likely stems from the inherent complexities of perfectly aligning machine-generated prosody and facial kinematics with human perceptual expectations, which are highly attuned to subtle inconsistencies.
Furthermore, the choice of generation parameters within the animation tools themselves warrants careful consideration. In this study, an "exaggerated mode" was utilized for the NVIDIA Audio2Face settings. While intended to ensure distinct emotional expression, this may have contributed to perceptions of unnatural or overly pronounced facial movements, as noted by some participants in the visual-only condition. Exploring "automatic" or more nuanced settings could yield more subtle and naturalistic expressions, potentially mitigating these issues and enhancing overall realism. Future work should systematically investigate the impact of these diverse settings. The inherent capabilities of technologies like MetaHuman, particularly features such as subsurface scattering for realistic skin rendering, offer significant potential in overcoming the uncanny valley. However, realizing this potential fully depends on harmonizing these advanced visuals with appropriately matched age-specific voice, precise lip synchronization, and nuanced emotional expressions driven by tools like Audio2Face.
5.1 Limitations
While this study provides key findings, certain limitations should be acknowledged:
  • Avatar Specificity: The findings are based on two specific MetaHuman avatars. Perceptions might differ with avatars of varying design, age representation, or art style.
  • Voice Realism and Appropriateness: The use of a young adult female voice for child avatars, though based on prior work, may have influenced perceptions of authenticity and emotional congruence. The availability and integration of high-quality, age-appropriate synthesized child voices remain a challenge.
  • Emotion Scope: The study focused on three basic emotions (joy, sadness, anger). The system’s efficacy in portraying a wider range or more complex emotional states (e.g., fear, surprise, confusion) was not assessed.
  • Participant Sample: Participants were recruited via an online platform (Prolific). While providing diversity, this sample may not fully represent the target end-users of specialized training systems, such as Child Protection Services (CPS).
  • Controlled Environment: The study was conducted using pre-rendered video clips. Interaction in a fully real-time, dynamic environment might elicit different perceptual responses.
5.2 Future Work
Several avenues for future research emerge from this study. The observation, echoed in user comments, that perceptions and responses may evolve with continued exposure highlights the need for longitudinal studies. Such research would offer valuable insights into adaptation, learning effects, and the long-term engagement potential of emotionally expressive avatars. Developing globally applicable systems necessitates exploring cultural variations in the expression and perception of emotions. Avatar design and emotional synthesis should account for these differences to ensure broader relatability and avoid misinterpretation. Finally, a critical next step is to conduct user studies with domain experts, such as professionals in CPS or law enforcement. Evaluating the effectiveness and usability of these emotionally expressive avatars within realistic child forensic interview training simulations will provide invaluable feedback. This will help identify areas for refinement to meet the specific, nuanced requirements of these high-stakes environments, ultimately assessing the practical impact of the system in its intended application context.
@@ 6 Conclusion
This research successfully establishes the feasibility of a novel, real-time architecture integrating Unreal Engine 5 MetaHuman and NVIDIA Omniverse Audio2Face for generating dynamic and nuanced facial emotions in AI-generated avatars. The core achievement lies in demonstrating this system’s capacity to produce recognizable emotional expressions, particularly for low-arousal states like joy and sadness, thereby marking a significant step towards creating more believable and effective virtual agents. These advancements are especially pertinent for high-stakes applications, such as virtual training simulations for child forensic interviews, where emotional realism can profoundly impact training efficacy.
Our findings underscore the differential perceptual impact of emotions based on arousal levels and the crucial role of multimodal congruence, revealing that while visual cues suffice for some emotions, accurately conveying high-arousal states like anger critically depends on synchronized audiovisual input. The study further pinpointed audiovisual synchronization and avatar morphology as pivotal factors influencing perceived realism and emotional clarity, offering key technical and design considerations for future development in the field. While challenges such as navigating the uncanny valley and refining the subtlety of expressions persist, this work clearly demonstrates the considerable potential of combining high-fidelity rendering with sophisticated audio-driven animation. Beyond specialized training, the principles and insights derived contribute to the broader endeavor of creating more naturalistic, engaging, and perceptually rich artificial characters across diverse interactive environments. This study provides a foundational contribution to the growing body of research on emotionally expressive avatars, paving the way for future innovations in human-AI interaction.

### v2

@@ 4.1.2 Emotion Recognition
term p.value OR OR_low OR_high
Emotion $\neq$ Intended 0.000 0.039 0.024 0.063
Audio+Visual 0.437 1.310 0.663 2.586
Avatar: Emory 0.002 0.202 0.073 0.562
Audio+Visual × Emotion $\neq$ Intended 0.976 1.010 0.547 1.864
Emotion $\neq$ Intended × Avatar: Emory 0.000 6.191 3.299 11.618
Audio+Visual × Avatar: Emory 0.000 4.045 1.931 8.472
3-way Interaction 0.000 0.212 0.091 0.493
Figure 6: Forest plot showing OR and 95% CI for predictors of emotion recognition accuracy. The dashed vertical line indicates no effect (OR = 1). Statistically significant predictors are clearly indicated. Reference levels: Avatar = Amelia, Condition = visual-only, Emotion = intended.
These findings reinforce the idea that while sadness can be conveyed effectively through visual cues alone, accurate recognition of emotions like joy often depends on audiovisual congruence, particularly for avatars with less prototypical or more ambiguous facial expressions [2].
Among all emotions, anger recognition was the most negatively affected in the visual-only condition. Ratings for Emory fell from $M=3.83$ to $M=1.43$, and for Amelia from $M=3.11$ to $M=2.17$. Silent anger was often misinterpreted as sadness, reflecting the high reliance of anger on auditory cues such as spectral roughness and vocal intensity [37, 38]. Visual signals alone, particularly AU4, were insufficient for conveying high-arousal affect.
Even in the audio+visual condition, Amelia’s anger ($M=3.11$) was frequently overshadowed by sadness ($M=3.71$), suggesting perceptual interference from her softer features and higher-pitched voice. Emory’s more angular structure and pronounced brow tension better conveyed anger ($M=3.83$ vs. $M=3.26$ for sadness), underscoring the role of morphology in affect salience. This may have been further influenced by his highly emotive voice, which could have reinforced the expression of anger by providing strong auditory cues.
@@
[…]
One plausible explanation for Amelia’s superior emotion recognition lies in her softer facial morphology, which aligns more closely with users’ expectations of emotional authenticity [39]. Her rounded features and gentle brow dynamics enhanced the perceived realism of low-arousal emotions like sadness and joy. In contrast, Emory’s angular face and more pronounced brow movements made high-arousal expressions like anger more salient [38, 37]. However, beyond facial structure, the alignment between voice characteristics and facial appearance may also have played a role. Amelia’s relatively soft and emotionally neutral voice appeared to match her visual design, potentially reinforcing the perception of coherence. In contrast, Emory’s highly expressive voice may have occasionally clashed with his facial expressions, leading to moments of perceptual mismatch and ambiguity. This highlights a key design trade-off: while subtle facial geometry can enhance believability for certain emotions, it may reduce clarity for others, especially when not fully synchronized with vocal affect.
4.1.3 Perceived Facial Expressions Realism
To investigate the perceived realism of facial expressions, we analyzed participants’ ratings across two experimental conditions (audio+visual and visual-only) and two avatar identities.
[…]
This enhancement likely stems from audiovisual interference. In the audio+visual condition, even slight mismatches in lip-sync or emotional prosody can disrupt the illusion of authenticity. The visual-only condition eliminates such dissonance, allowing facial cues to be evaluated in isolation, which may reduce perceptual scrutiny and heighten tolerance for animation artifacts.
[…]
Figure 7: Mean ratings ($M\pm SD$) perceived Facial Expressions realism for each video clips under both audio+visual and visual-only conditions.
[…]
term p.value OR OR.lower OR.upper
Visual-only 0.526 1.568 0.391 6.294
Avatar: Emory 0.001 0.211 0.084 0.528
Scenario: Joy 0.722 0.850 0.348 2.077
Scenario: Sad 0.228 0.575 0.234 1.413
Visual-only × Avatar: Emory 0.055 3.584 0.972 13.208
Visual-only × Scenario: Joy 0.198 2.387 0.635 8.981
Visual-only × Scenario: Sad 0.023 4.554 1.232 16.829
Avatar: Emory × Scenario: Joy 0.216 2.254 0.622 8.165
Avatar: Emory × Scenario: Sad 0.875 1.108 0.310 3.963
Visual-only × Avatar: Emory × Scenario: Joy 0.081 0.188 0.029 1.225
Visual-only × Avatar: Emory × Scenario: Sad 0.115 0.229 0.037 1.433
Figure 8: Forest plot showing OR with 95% CI for the model predictors of perceived realism of facial expressions. The vertical dashed line indicates no effect (OR = 1). Statistically significant predictors are clearly indicated. Reference levels: Avatar = Amelia, Condition = Audio+Visual, Scenario = Angry.
[…]
In summary, low-arousal emotions like sadness appear more effectively conveyed through visuals alone, as unsynchronized audio can introduce perceptual dissonance that reduces realism. These findings underscore the importance of condition-sensitive avatar design, especially in emotionally nuanced applications such as child-simulation training, where voice-age congruence and expressive fidelity are crucial to achieving believable interaction.
Figure 9: Mean ratings ($M\pm SD$) of perceived realism across different aspects (facial expressions, visual appearance, voice tone, dialogue content) and empathy, compared between audio+visual and visual-only conditions.
[…]
term p.value OR OR_low OR_high
Realism Audio+Visual vs. Visual-Only 0.038 0.258 0.072 0.925
Visual Appearance vs. Facial Expressions 0.765 0.865 0.333 2.246
Interaction: Audio+Visual × Visual Appearance 0.587 1.434 0.391 5.264
Empathy Audio+Visual vs. Visual-Only 0.397 0.693 0.295 1.614
4.1.4 Perceived Realism and Empathy
The results summarized in Table 7 and Figure 9 illustrate how multimodal cues (audio and visual) influenced participants’ perception of realism and empathy towards emotionally expressive avatars.
In the audio+visual condition, auditory elements played a major role in shaping perceptions. Voice tone emerged as the strongest contributor, with an average realism rating of $M=4.00$, followed by dialogue content ($M=3.86$). However, in the visual-only condition, visual components, particularly facial expression realism and visual appearance, received higher ratings ($M=3.86$ and $M=3.80$, respectively) compared to the audio+visual condition ($M=3.29$ and $M=3.40$, respectively). This difference may partly reflect the fact that participants in the visual-only condition evaluated fewer realism contributors, with their attention concentrated on the available visual aspects. In contrast, participants in the audio+visual condition distributed their judgments across four factors, including voice tone and dialogue content, which may have diluted individual ratings for visual features.
[…]
Collectively, these findings highlight a critical trade-off in avatar design: while audio cues enhance realism for high-arousal emotions, subtle audiovisual mismatches can negatively impact perceived authenticity. Under the conditions examined, removing audio increased visual credibility without substantially affecting emotional engagement.
@@ 4.2 Qualitative Feedback
Participants provided reflective qualitative feedback on what stood out the most about the avatars. These reflections revealed distinct and important observations regarding emotional expressiveness, realism, and immersion, providing directions for enhancing avatar technologies in interactive simulations. Two independent coders analyzed the responses. Discrepancies in coding were discussed and resolved collaboratively to ensure consistency and analytical rigor.
@@ 5 Discussion
This study offers several important lessons regarding the perceptual and design challenges of building emotionally expressive child avatars. Rather than emphasizing the system’s technical capabilities, our findings point to deeper tensions in achieving believable emotional portrayals through multimodal integration. One key insight is that perceived realism depends not on modality fidelity in isolation, but on the coherence across modalities. Moreover, participant responses varied considerably, revealing the inherently subjective nature of emotional perception. What one viewer perceives as lifelike or relatable, another may find unsettling or unnatural. This variability underscores the importance of designing emotionally expressive avatars that can accommodate a range of user sensitivities. When auditory and visual channels are not well aligned, user experience suffers. In some cases, visual-only presentations were even rated as more realistic than audio-visual ones. Silencing the audio removes a major source of incongruence, but it can also weaken the communication of certain emotions, particularly those that involve high arousal.
The study highlights a recurring trade-off. Auditory cues are necessary for disambiguating intense emotional states such as anger, yet their inclusion risks drawing attention to imperfections in lip-sync or mismatched vocal tone. Synchronizing prosody with facial animation is not a small technical issue. It plays a key role in making avatars believable and emotionally clear. Participant feedback supports this, with several comments describing distracting or unsettling effects when lip movements did not match the dialogue or when the voice tone conflicted with the avatar’s appearance.
A second lesson concerns the role of avatar morphology in shaping emotional interpretation. The "Amelia" avatar, with softer facial features, was generally perceived as more believable in low-arousal contexts such as sadness or joy. In contrast, the more angular "Emory" avatar was more effective in conveying anger. These outcomes suggest that certain facial structures are better suited to subtle, empathic expressions, while others may more effectively convey direct, high-intensity emotions. Designers of virtual agents may benefit from selecting or customizing facial geometry based on the emotional demands of the specific use case.
Additionally, the study raises questions about the optimal configuration of animation parameters. Our use of the “exaggerated” mode in Audio2Face, intended to increase clarity, some participants in the visual-only condition perceived the resulting facial movements as overly intense or unnatural. This suggests a potential trade-off between expressiveness and realism, especially in training contexts where exaggerated expressions may break immersion or reduce credibility. Exploring "automatic" or more customized settings could help address this issue.
Overall, this work reinforces that perceived believability is not simply a function of system complexity or realism in individual channels, but rather the result of coherent integration across modalities and sensitivity to user perception. By focusing on these lessons, future systems can be better tailored to the emotional and communicative needs of users, especially in sensitive and high-stakes environments.
5.1 Limitations and Future Work
While this study offers several valuable insights into the design and perception of emotionally expressive child avatars, it also presents a number of limitations that inform directions for future research. The findings are based on two specific MetaHuman avatars, and user perceptions may differ with avatars of varying visual design, age representation, or artistic style. Moreover, although prior work supported the use of a young adult female voice for child avatars, this choice may have influenced participants’ perceptions of emotional authenticity and audiovisual coherence. Integrating high-quality, age-appropriate synthesized voices that align more naturally with the avatars’ appearance and emotional tone remains a significant technical and perceptual challenge.
The emotional scope of the study was also limited to three basic emotions: joy, sadness, and anger. As a result, the system’s ability to portray a broader range of complex or subtle emotional states (such as fear, surprise, or confusion) remains untested. Additionally, participant recruitment via the Prolific platform, while offering demographic diversity, may not fully capture the perspectives of domain-specific end-users such as Child Protection Services (CPS) professionals. The controlled, pre-rendered nature of the video stimuli further limits ecological validity. Responses observed in dynamic, real-time interactions may differ substantially due to the role of responsiveness and timing in social perception.
These limitations point to several promising avenues for future work. Longitudinal studies are needed to explore how user perceptions evolve with repeated or sustained exposure to emotionally expressive avatars, potentially revealing learning effects and long-term engagement dynamics. Future research should also expand the emotional repertoire of avatars to better reflect the range of expressions relevant in sensitive contexts such as forensic interviews. Critically, upcoming evaluations should involve domain experts such as CPS or law enforcement professionals to assess the system’s usability, realism, and training efficacy in practical, high-stakes environments. Such studies will help refine both the technical and perceptual aspects of the system, ensuring its relevance and effectiveness in its intended application context.
@@ 6 Conclusion
This study highlights a central challenge in the development of emotionally expressive child avatars: effective multimodal integration remains difficult to achieve. While our system demonstrates the technical feasibility of real-time emotional animation using Unreal Engine 5 MetaHuman and NVIDIA Audio2Face, the results reveal that audiovisual misalignment can undermine perceived authenticity. Notably, participants often rated visual-only clips as more realistic than audio-visual ones, suggesting that incomplete or poorly matched modalities may detract from overall believability.
This is especially relevant in sensitive applications such as child forensic interview training, where avatars must evoke trust and empathy. Our findings show that high-arousal emotions like anger depend heavily on synchronized vocal and facial cues, while lower-arousal states like sadness and joy can be conveyed more effectively through visual expressions alone. Additionally, factors such as avatar morphology and voice-age congruence significantly shaped emotional perception.
Rather than pursuing realism within individual modalities, this study underscores the importance of holistic and context-sensitive design. Emotional credibility arises only when voice, appearance, and facial behavior are cohesively aligned. These insights offer practical guidance for improving the design of virtual agents in emotionally demanding settings. Ultimately, consistent integration across modalities, rather than fidelity in isolation, is essential for achieving believable human–AI interaction.
Appendix: Supplementary Material
This appendix includes supplementary statistical tables (Table 8 and Table 9) that report model comparisons and detailed regression results supporting the analyses presented in Section 4.
[…]
Variable
Mean
visual-only
Mean
audio+visual
p-value Cohen’s d
FDR-corrected
p-value
Facial exp. 3.429 2.800 0.006 0.680 0.021
Angry 1.829 2.571 0.019 -0.575 0.053
Sad 3.343 3.629 0.360 -0.221 0.520
Joy 1.429 1.171 0.119 0.378 0.222
Fear 2.229 1.857 0.192 0.315 0.312
Emory_Sad Disgust 1.971 1.971 1.000 0.000 1.000
Facial exp. 3.514 2.914 0.028 0.535 0.074
Angry 1.429 3.829 $0$2.565 0.000
Sad 2.000 3.257 $0$0.985 0.001
Joy 2.143 1.086 0.000 1.329 0.000
Fear 1.714 1.914 0.448 -0.182 0.540
Emory_Angry Disgust 1.486 2.743 $0$1.029 0.001
Facial exp. 3.514 3.171 0.226 0.292 0.339
Angry 1.400 1.057 0.008 0.661 0.027
Sad 1.743 1.114 0.005 0.717 0.021
Joy 2.914 3.771 0.005 -0.690 0.021
Fear 1.571 1.229 0.082 0.423 0.177
Emory_Joy Disgust 1.743 1.171 0.009 0.656 0.027
Facial exp. 4.000 3.343 0.006 0.678 0.021
Angry 2.314 2.571 0.425 -0.192 0.540
Sad 4.086 4.114 0.901 -0.030 0.925
Joy 1.200 1.086 0.224 0.293 0.339
Fear 2.314 2.086 0.455 0.179 0.540
Amelia_Sad Disgust 1.800 2.429 0.036 -0.513 0.081
Facial exp. 3.629 3.543 0.726 0.084 0.787
Angry 2.171 3.114 0.004 -0.714 0.021
Sad 2.371 3.714 $0$0.997 0.001
Joy 1.829 1.143 0.005 0.700 0.021
Fear 1.857 2.000 0.589 -0.130 0.656
Amelia_Angry Disgust 1.943 2.114 0.577 -0.134 0.656
Facial exp. 3.914 3.486 0.098 0.401 0.200
Angry 1.086 1.000 0.179 0.325 0.304
Sad 1.143 1.057 0.400 0.203 0.540
Joy 4.371 4.200 0.405 0.200 0.540
Fear 1.143 1.114 0.754 0.075 0.795
Amelia_Joy Disgust 1.229 1.057 0.119 0.377 0.222
[…]
Audio+Visual Visual-Only
Variable p-value
Mean
Amelia
Mean
Emory
Cohen’s d
FDR-corrected
p-value
p-value
Mean
Amelia
Mean
Emory
Cohen’s d
FDR-corrected
p-value
Facial exp. 0.001 3.343 2.800 0.591 0.010 0.001 4.000 3.429 0.646 0.005
Angry 1.000 2.571 2.571 0.000 1.000 0.088 2.314 1.829 0.297 0.143
Sad 0.020 4.114 3.629 0.414 0.045 0.018 4.086 3.343 0.419 0.041
Joy 0.263 1.086 1.171 -0.192 0.338 0.118 1.200 1.429 -0.271 0.163
Fear 0.244 2.086 1.857 0.201 0.338 0.619 2.314 2.229 0.085 0.619
Sad Disgust 0.014 2.429 1.971 0.440 0.041 0.280 1.800 1.971 -0.186 0.336
Facial exp. 0.002 3.543 2.914 0.578 0.010 0.579 3.629 3.514 0.095 0.613
Angry 0.002 3.114 3.829 -0.582 0.010 0.006 2.171 1.429 0.495 0.016
Sad 0.009 3.714 3.257 0.466 0.041 0.113 2.371 2.000 0.275 0.163
Joy 0.571 1.143 1.086 0.097 0.643 0.162 1.829 2.143 -0.242 0.208
Fear 0.609 2.000 1.914 0.087 0.644 0.483 1.857 1.714 0.120 0.543
Angry Disgust 0.012 2.114 2.743 -0.451 0.041 0.004 1.943 1.486 0.516 0.016
Facial exp. 0.054 3.486 3.171 0.337 0.108 0.041 3.914 3.514 0.358 0.075
Angry 0.160 1.000 1.057 -0.243 0.262 0.006 1.086 1.400 -0.498 0.016
Sad 0.324 1.057 1.114 -0.169 0.389 0.002 1.143 1.743 -0.565 0.012
Joy 0.020 4.200 3.771 0.413 0.045 0.000 4.371 2.914 1.025 0.000
Fear 0.254 1.114 1.229 -0.196 0.338 0.026 1.143 1.571 -0.392 0.053
Joy Disgust 0.103 1.057 1.171 -0.283 0.186 0.006 1.229 1.743 -0.495 0.016
[…]
Conversely, in scenarios depicting high-arousal emotions (Emory-Angry, Amelia-Angry, and Emory-Joy), the presence of audio improved emotion recognition and overall expressiveness. This pattern indicates that synchronized vocal cues can enhance the clarity of emotional intent, particularly when facial expressions alone provide insufficient affective information.
Table 9 provides a direct comparison between the two avatars. Across both conditions, Amelia was consistently rated as more realistic, with higher accuracy in emotion identification for sadness and joy. Notably, in the audio-visual condition, Emory was rated more believable in the angry scenario, suggesting that his angular features may better support high-arousal affect when complemented by vocal intensity. In contrast, Amelia’s rounded, soft-featured face reliably conveyed emotional states across modalities, enabling consistent perception even without audio.
Taken together, these two tables reinforce our central claim: perceived believability depends critically on the interplay between audiovisual congruence and facial geometry.